# NEDS-SLAM: A Novel Neural Explicit Dense Semantic SLAM Framework using 3D Gaussian Splatting

### v2

@@ V Conclusion and Limitations
The proposed NEDS-SLAM is an end-to-end semantic SLAM system based on 3D Gaussian splatting. By integrating a Spatially Consistent feature fusion model, NEDS-SLAM effectively addresses the challenges of robustly estimating semantic labels with pre-trained models, significantly enhancing semantic reconstruction performance The proposed Virtual View Pruning method leverages the capabilities of differentiable Gaussian splatting for fast and realistic novel view synthesis. It effectively eliminates outlier GS points in the SLAM process, significantly enhancing the reconstruction quality of the displayed neural radiance fields.
The experiment using public datasets further validated the effectiveness of NEDS-SLAM. However, we also identified existing shortcomings during the experiment. The Virtual view pruning method, which generates virtual views in the keyframe reconstruction pipeline, increases computational load and can affect real-time performance. In the future, we plan to further optimize the virtual view method and take into account semantic reconstruction in dynamic scenes.

### v3

@@ V Conclusion and Limitations
The proposed NEDS-SLAM is an end-to-end semantic SLAM system based on 3DGS. By integrating a Spatially Consistent feature fusion model, NEDS-SLAM effectively addresses the challenges of robustly estimating semantic labels with pre-trained models, significantly enhancing semantic reconstruction performance. The Virtual Camera View Pruning method uses differentiable Gaussian splatting for quick and realistic novel view synthesis. It removes outlier gaussians during SLAM, significantly improving the reconstruction quality of neural radiance fields.
The experiment with public datasets confirmed NEDS-SLAM’s effectiveness but revealed some shortcomings. The virtual camera view pruning method improves mapping speed by removing more gaussians. However, increasing the frequency of VCVP usage also raises computational load, indicating room for further optimization. Future plans include optimizing and incorporating semantic reconstruction for dynamic scenes.